# Auto Arena of LLMs: Automating LLM Evaluations with Agent Peer-battles and Committee Discussions

### v1

@@ 1 Introduction
Since ChatGPT and GPT-4 [23] gained popularity, large language models (LLMs) have risen to the forefront of technological innovation, capturing broad industry and social interests. This enthusiasm has spurred numerous organizations to release their own LLMs [27, 26]. However, the rapid pace at which these models are released and updated poses a significant challenge for users attempting to understand their capabilities and monitor their evolution. Consequently, there has been a pressing demand for comprehensively evaluating LLMs recently.
@@
One line of research conducts automatic evaluation with static datasets. Among these, static datasets with predefined metrics, such as GSM8k [9] and MMLU [15], are constructed with aspect-specific input-output pairs, such as questions and their corresponding answers. Given the questions, the LLM-produced answers are compared to ground-truth answers using metrics such as accuracy. However, these approaches suffer significantly from contamination concerns [24], where models may have been inadvertently exposed to elements of the test datasets during training, thereby skewing evaluation results. The rigid ground-truth answers also limit their utility in assessing models’ performance on general or open-ended questions. Model-based evaluation, such as MT-Bench [32] and AlpacaEval [12], provides an alternative for evaluating LLMs on open-ended questions. These methods typically ask two models to generate responses to the same open-ended question and then employ a strong judge model (e.g., GPT-4) to choose the better response. However, the static question sets still pose the issue of contamination. Additionally, the assumption of the existence of a strong judge model makes the evaluation framework less generalizable.
@@
Aside from automated evaluations, human assessment, although requiring significant manual efforts, remains the gold standard for most users. A noticeable example is Chatbot Arena [32], which is a crowdsourced voting platform that gathers anonymous votes on LLM performances and calculates ELO scores to rank these models. The resulting leaderboard222  https://leaderboard.lmsys.org/ is widely considered a trustworthy indicator of an LLM’s general capabilities. However, a reliable model evaluation on this platform must be supported by more than 10k human votes, which requires considerable time and effort. Consequently, when newly developed models enter the scene, they often struggle to quickly amass a significant number of votes. Moreover, this strong reliance on human votes limits its application in various scenarios. For example, the performance of non-English languages is difficult to estimate, as most of the queries on the platform are in English. The completely open participation may also result in uneven evaluation quality.
Figure 1: An illustration of Auto Arena of LLMs.
@@
To provide automatic, reliable, and human-like LLM evaluations, we propose Auto Arena of LLMs (Auto-Arena), a framework that automates the whole LLM evaluation process with LLM agents. As shown in Figure 1, the Auto-Arena framework consists of three stages. Firstly, an LLM examiner agent is tasked with generating questions, mimicking real-life users inputting queries. Secondly, two candidate LLMs interact with each other and engage in a multi-round peer battle by answering the seed question individually, criticizing the opponent’s weaknesses, and raising targeted follow-up queries to challenge the opponent further. During the multi-round battle process, the LLM’s true capabilities are drawn out and performance gaps become visible. Lastly, a committee of LLM judge agents collectively discusses and evaluates the ability of the two candidates, mimicking the human voting process. By automating the entire evaluation process with LLM agents, we alleviate data contamination concerns in static datasets, reduce single-model bias with collective decision-making, and avoid long wait times for new models entering the human voting platform.
@@
To verify the evaluation framework, we run an extensive experiment with 17 models. Compared to static and model-based benchmarks, Auto-Arena increases the Spearman correlation with human preferences by 4.5%, resulting in state-of-the-art alignment. Before and after peer battles, the Spearman correlation with human preferences increases by 46.4%, verifying our hypothesis the peer battles can better display performance gaps. Before and after committee discussions, committee agreement increases by 20%, showcasing the effectiveness of the committee discussion mechanism. By studying the peer battles, we discover intriguing LLM agent behaviors such as self-improvement and competitive actions. Using Chinese as an example, we also show that such an automated framework is easily extendable to non-mainstream languages and domains. Along with the project, we release a leaderboard, a website with demos, and the codebase. We are actively maintaining the leaderboard and altering the questions to provide reliable results in a timely manner.
@@ 2 Related Work
Since LLMs demonstrated general and zero-shot capabilities, researchers have been attempting to use one single strong LLM to judge other LLMs’ outputs. LLM-as-a-judge [32] introduces Chatbot Arena and MT-Bench, a multi-turn question set to mimic real-life user queries. Using a strong LLM judge, such as GPT-4, to evaluate the MT-Bench results achieves over 80% agreement with human preferences, showcasing the potential of such methods. Besides such usefulness evaluations, some attempt to use LLM-as-a-judge to examine the knowledge boundaries of LLMs. Language-Model-as-an-Examiner [2] constructs the LMExamQA dataset, which comprises of LM-generated, knowledge-intensive questions covering three cognitive levels, including knowledge memorization, knowledge comprehension, and knowledge analysis. These questions are required to be answered by the LLM itself correctly, limiting their complexity. Then, the LM evaluator and the candidate interact in a series of follow-up queries and the evaluator rates the responses on dimensions including accuracy and factuality. Multiple evaluators examine the model independently and vote for a final result. KIEval [29] also incorporates an LLM-powered “interactor” role to examine deep comprehension of knowledge, which is shown to mitigate contamination issues. By utilizing the LLMs to generate questions, these methods can effectively mitigate the contamination concerns on static datasets. However, such evaluations only examine the LLMs’ capabilities on answering knowledge-intensive, closed-domain, and simpler questions, as they rely on the examiner’s internal knowledge. They also require the examiner to interact with each candidate individually, creating computational overheads and limiting the scope of queries.
@@
Moreover, using a single LLM judge still suffer from a bias towards LLM-generated summaries [20], inflated scores in multilingual evaluation [14], verbosity bias [12], and difficulties when evaluating candidates with close performance [25]. To mitigate these problems, some recent works propose using multiple agents for collaborative evaluation, mimicking a peer review process. Some works simulate different personas with the same LLM. DRPE [28] uses multi-roleplayer prompting to mimic different roles with the same base model and integrate multiple outputs as votes for the final results. ChatEval [6] simulates different evaluator personas with the same base model to engage in debates, reaching a final evaluation result. While these methods come from the same motivation as our committee discussion component, they are still built on existing static datasets and do not mitigate the model-specific bias as they still use the same LLM to simulate different personas. To further incorporate different LLMs as evaluators, PRD [17] allows 2 LLMs to discuss an evaluation and assigns higher voting weights to the LLM reviewers with stronger capabilities. Peer-review-in-LLMs [22] further optimizes the voting weights as a learnable parameter. WideDeep [30] sets up a multi-layer neural network structure where each LLM evaluator functions as neurons. PRE [8] selects a small group of reviewers to produce evaluations individually and then asks a “chair” to produce the final evaluation using the aggregated evaluations. They show that the multi-agent approach effectively mitigates single-model bias. This line of work is similar to our “LLM judge committee”. However, they are still limited on static datasets and specific domains. Moreover, they lack the collective discussion component, which is shown to improve agreement significantly.
@@
Compared to using LLMs for evaluations, LLM-debate is a relatively new research area. Cohen et al. [10] shows that LLM interactions and cross-examination can effectively discover factual errors. Debate [11] shows that multi-agent debate can improve factuality and reasoning abilities. MAD [19] shows that LLM debate can encourage divergent thinking, which is helpful for tasks that require deep levels of contemplation. Khan et al. [16] shows that even non-expert weak LLMs can supervise expert LLMs if we allow the two LLM experts to engage in debates. Moreover, Zhao et al. [31] and Gu et al. [13] show interesting case studies where LLMs are engaged in simulated competitive environments, where they demonstrate human-like strategies.
@@ 3 The Auto Arena Framework
The Auto-Arena framework consists of three stages: question generation, peer battles, and committee discussions. These three stages are run sequentially and fully automated with LLM-based agents. The overall design of our Auto-Arena framework is presented in Figure 1. All prompts are included in Appendix A.1.
@@ 3.1 Question Generation
For debate questions, as using a static dataset could incur data contamination concerns and result in unfair evaluations, we ask an LLM examiner agent to dynamically generate questions. The examiner agent could be any capable LLM. As the question simply serves as a starting point for the peer battle, which is the primary focus of evaluation, the question quality has relatively less impact on the overall evaluation framework. Similar to MT-Bench [32], we utilize 8 common categories in real-life conversations: writing, roleplay, extraction, reasoning, math, coding, STEM knowledge, and humanities/social science knowledge. In the prompt, the examiner is provided with a sample question and encouraged to generate diverse and difficult questions to ensure the depth and width of the debates examined. A subset of the generated questions is shown in Appendix A.2.
Specifically, as the examiner agent will also participate in the following debates, previous methods [2] could incur self-enhancement bias as the examiner agents only devise questions that they are confident about. We try to avoid such bias with the following two designs: 1. We do not disclose to the examiner that it will participate in this tournament. 2. We do not ask the examiner to only generate questions that it can solve. Thus, the resulting questions should not cause much enhancement bias for the examiner agent itself.
@@ 3.2 Peer Debate
Figure 2: The process of a Lincoln-Douglas-style peer-battle with the actions used.
After the questions are drafted, we conduct peer battles around these questions among the candidate LLMs. In one peer battle, two candidate LLMs debate around the given question, point out the opponent’s weaknesses, and raise follow-up questions that the opponent may fail to respond to.
In the peer battle, each candidate LLM can utilize 4 types of actions: 1. Think, where the candidate thinks about the question and plans its strategy. This action is hidden to the opponent at all times. 2. Respond, where the candidate responds to the aforementioned question. 3. Criticize, where the candidate points out the loopholes and mistakes in the opponent’s previous responses. 4. Raise, where the candidate raises follow-up questions that are specifically designed to make the opponent expose its weaknesses.
The peer battle is designed according to the Lincoln-Douglas debate format  [^3] https://en.wikipedia.org/wiki/Lincoln-Douglas_debate_format, which is a competitive one-to-one debate used for popular competitions such as National Speech and Debate Association competitions. Overall, the peer battle consists of 3 rounds, where the candidates take turns to speak. The entire dialogue history is visible to both candidates. The process is illustrated in Figure 2. In the first round, A gives an initial response to the examiner’s question; B criticizes the weaknesses in A’s response and raises a targeted follow-up question; and A responds to B’s question. In the second round, A and B are reversed: B gives an initial response to the examiner’s question (without seeing A’s response); A criticizes and raises questions; and B responds to A’s question. In the third round, A and B cross-examine each other. A starts by criticizing B’s previous loopholes and raises follow-up questions. After responding, B also criticizes A’s loopholes and raises questions. A concludes the battle by responding again. In this process, both A and B get an equal number of each action to ensure fairness. To further reduce position bias, A and B’s order is randomly shuffled at the beginning of each debate.
Depending on which turn it is, we provide an action guide to the candidate, specifying the objectives and corresponding actions for this turn. Similar to human debate competitions, we time the candidates by providing a maximum length, which is also specified in the prompts. Any responses beyond the required length will be cut off. This design also mitigates verbosity bias in LLM-as-a-judge [32], where LLM judges prefer longer and more verbose responses.
@@ 3.3 Committee Discussions
After the peer battle takes place, a committee of LLM judge agents will collectively determine the winner. The committee is always selected as the best five LLMs according to the current ranking. In the first round, the committee is initialized with MMLU [15] scores to approximate LLM performances. They will first be asked to read through the battle history, elaborate judgment reasons, and give a verdict on whether A is better, or B is better, or if there is a tie.
After the initial judgments are formed, the committee engages in a discussion. Each judge will read the other judge’s verdicts and decide whether to adjust the ratings and elaborate on the reasons. The collective intelligence introduces diverse viewpoints, improves judgment quality, and mitigates single-model bias. Finally, the winning candidate is decided by the majority voting of the discussed judgments.
For logical-reasoning questions that have ground-truth answers (reasoning, code, math), LLM-as-a-judge is known to show weak performances in judging the quality of responses. We adopt prior approaches to establish the reference-based judge [32]. Specifically, we utilize the strongest model (according to the current ranking) to generate a reference answer and provide it to the judge when evaluating the peer battle.
@@ 4.1 Experimental Setup
For the main experiment, we first select 7 models from the top 30 list on Chatbot Arena with more than 10k votes each. The models are selected to be the best or the latest models from each popular model family at the time of experiments. To construct a comprehensive leaderboard [^4] https://huggingface.co/spaces/Auto-Arena/Leaderboard, we further consider 10 more models.
@@
For the baselines, we consider widely used evaluation benchmarks, including static datasets with fixed metrics and model-based metrics.
  1. 1.
    Static datasets with fixed metrics
    1. (a)
      MMLU (Massive Multitask Language Understanding) [15], an extensive benchmark dataset that covers 57 subjects and tests both world knowledge and problem-solving ability.
    2. (b)
      OpenLLM Leaderboard [3], a leaderboard calculated by averaging performance metrics on 6 key benchmarks, including MMLU.
  2. 2.
    Static datasets with model-based metrics
    1. (a)
      MT-Bench [32], a multi-turn question set consisting of 80 questions. Model responses are graded by GPT-4.
    2. (b)
      MT-Bench Hard [18], a benchmark dataset with 1,000 challenging user queries collected on Chatbot Arena. Model responses are graded by GPT-4.
    3. (c)
      Length-controlled AlpacaEval [12], an automatic benchmark based on the AlpacaFarm evaluation set, which tests the ability of models to follow general user instructions. Model responses are graded by GPT-4.
@@
Among the 7 participants, we conduct a swiss-style tournament: For $n$ participants, instead of pairing each participant with $(n-1)$ others, a swiss-tournament pairs each player with $log_{2}(n)$ players of similar rankings without repeats. This design effectively reduces computational costs from $O(n^{2})$ to $O(nlog_{2}(n))$.
Each pair of candidates engage in 40 peer battles, with 5 questions from each of the 8 categories. The questions are generated by GPT-4. As each battle consists of 3 rounds (each candidate speaks for 4 times), we expect the competition scale to be approximately the same as MT-Bench (80 questions, each candidate speaks twice).
In the tournament, the rating scores are provided by calculating the ELO rating system [1, 4], which has become the standard practice in competitive games such as chess. Similar to the Chatbot Arena score calculation procedure [7], we compute the Bradley-Terry (BT) coefficients [5] for better statistical estimation.
We initialize the Swiss tournament rankings according to MMLU scores, which is a static approximate of model performances. At the end of each pairing, we re-calculate ELO scores of current models. The committee is selected to be the remaining 5 models (besides the candidates). When the participant number increases, the committee is selected as the best 5 LLMs based on current ELO rankings. After the initial judgments, the committee members engage in one round of discussion.

### v2

@@ 1 Introduction
Since ChatGPT and GPT-4 [34] gained popularity, Large Language Models (LLMs) have risen to the forefront of technological innovation, capturing broad industry and social interests [40]. This enthusiasm has spurred numerous organizations to release their own LLMs [38, 37]. However, the rapid pace at which these models are released and updated poses a significant challenge for users attempting to understand their capabilities and monitor their evolution. Consequently, there has been a pressing demand for comprehensively evaluating LLMs recently [10].
@@
One line of these evaluation researches conducts automatic evaluation with static datasets. Among these, static datasets with predefined metrics, such as GSM8k [13] and MMLU [22], are constructed with aspect-specific input-output pairs, such as human exam-type questions and their corresponding answers. Given the questions, the LLM-produced answers are compared to ground-truth answers using metrics such as accuracy. However, these approaches suffer significantly from contamination concerns [35], where models may have been inadvertently exposed to elements of the test datasets during training, thereby skewing the evaluation results. The closed-formed ground-truth answers also limit their utility in assessing models’ performances on general or open-ended questions. Model-based evaluation, such as MT-Bench [44] and AlpacaEval [19], provides an alternative for evaluating LLMs on open-ended questions. These methods typically ask two models to generate responses to the same open-ended question and then employ a strong judge model (e.g., GPT-4) to choose the better response. However, the static question sets still pose the issue of contamination. Additionally, the assumption of the existence of a strong judge model makes the evaluation framework less generalizable.
@@
Aside from automated evaluations, human assessment, although requiring significant manual efforts, remains the gold standard for most users. A noticeable example is Chatbot Arena [44], a crowdsourced voting platform that gathers anonymous votes on LLM performances and calculates ELO scores to rank these models. The resulting leaderboard [^2] https://leaderboard.lmsys.org/ is widely considered a trustworthy indicator of an LLM’s general capabilities. However, a reliable model evaluation on this platform must be supported by a large number of human votes, which requires considerable time and effort. Consequently, when newly developed models enter the scene, they often struggle to quickly amass a large number of votes. Moreover, this strong reliance on human votes limits its application in various scenarios. For example, the performance of non-English languages is difficult to estimate, as most queries on the platform are in English. The completely open participation may also result in uneven evaluation quality.
Figure 1: An illustration of Auto Arena of LLMs (Auto-Arena).
@@
To provide automatic, reliable, and human-like LLM evaluations, we propose Auto Arena of LLMs (Auto-Arena), a framework that automates the entire LLM evaluation process with LLM agents. As shown in Figure 1, the Auto-Arena framework consists of three stages. Firstly, an LLM examiner agent is tasked with generating questions, mimicking real-life users posting queries. Secondly, two candidate LLMs interact with each other and engage in a multi-round peer-battle by answering the seed question individually, criticizing the opponent’s weaknesses, and raising targeted follow-up queries to challenge the opponent further. During the multi-round battle process, the LLM’s true capabilities are drawn out and performance gaps become more visible. Lastly, a committee of LLM judge agents collectively discusses and evaluates the ability of the two candidates, mimicking the human voting process. By automating the entire evaluation process with LLM agents, we alleviate data contamination concerns in static datasets, reduce single-model bias with collective decision-making, and avoid long wait times for new models entering the human voting platform.
@@
To verify the evaluation framework, we run an extensive experiment with 17 LLMs. Compared to static and model-based benchmarks, Auto-Arena increases the Spearman correlation with human preferences by 4.5%, resulting in state-of-the-art alignment. Before and after peer battles, the Spearman correlation with human preferences increases by 46.4%, verifying our hypothesis that the peer battles can better display performance gaps. Before and after committee discussions, committee agreement increases by 20%, showcasing the effectiveness of the committee discussion mechanism. By studying the peer battles, we also discover intriguing LLM agent behaviors such as self-improvement and competitive actions. Using Chinese as an example, we further show that such an automated framework is easily extendable to non-mainstream languages and domains. Along with the project, we release a leaderboard, a website with demos, and the codebase. We are actively maintaining the leaderboard and altering the questions to provide reliable results in a timely manner.
@@ 2 Related Work
As LLMs evolve quickly, deriving trustworthy evaluations of their capabilities has become a challenge. Current evaluation methods can be divided into manual evaluation [5] and automatic evaluations. We primarily focus on automatic evaluations as they deliver more timely feedback. Automatic evaluations mainly consist of static datasets with predefined metrics and model-based metrics. Static datasets with predefined metrics [22, 6] suffer from contamination concerns and inflexibility in evaluating open-ended responses. Among model-based metrics, MT-Bench [44] and AlpacaEval [19] use LLM-as-a-judge to ask GPT-4 to compare model responses to a static dataset of questions. The model’s judgments achieve over 80% agreement with human preferences, proving the usability of using LLMs to evaluate response quality. Language-Model-as-an-Examiner [5] asks an LM examiner to construct knowledge-intensive questions within its memory, interact with the candidate in a series of follow-up queries, and rate the responses on dimensions including accuracy and factuality. KIEval [41] also incorporates an LLM-powered “interactor” role to examine deep comprehension of knowledge, which is shown to mitigate contamination issues on static datasets. However, such single-judge evaluations require the examiner to interact with each candidate parallelly, creating computational overheads and limiting the scope of queries. They also suffer from single-model bias, including bias towards LLM-generated summaries [28], inflated scores in multilingual evaluation [21], verbosity bias [19], and difficulties when evaluating candidates with close performance [36]. Therefore, some attempt to employ multiple judges agents. DRPE [39] uses multi-roleplayer prompting to mimic different roles with the same LLM and integrate outputs as votes for the final results. ChatEval [9] simulates different evaluator personas with the same base model to engage in debates, reaching a final evaluation result. While these methods come from the same motivation as our committee discussion component, they do not mitigate the model-specific bias as they use the same base LLM to simulate different personas. To further incorporate different LLMs as evaluators, PRD [25] allows 2 LLMs to discuss an evaluation and assigns higher voting weights to the LLM reviewers with stronger capabilities. Peer-review-in-LLMs [31] optimizes the voting weights as a learnable parameter. WideDeep [42] sets up a multi-layer neural network structure where each LLM evaluator functions as neurons. PRE [12] selects a small group of reviewers to produce evaluations individually and then asks a “chair” to produce the final evaluation using the aggregated evaluations. They show that the multi-agent approach effectively mitigates single-model bias. This line of work is similar to our “LLM judge committee” component. However, they are still limited to static datasets and specific domains.
@@
Besides LLM evaluations, some works study competitive behaviors in LLMs, which is relevant to the LLM peer-battles in Auto-Arena. LM vs LM [14] shows that LLM cross-examinations can effectively discover factual errors. Debate [18] shows that multi-agent debate can improve factuality and reasoning. In MAD [27], LLM-debate can encourage divergent thinking, which helps tasks that require deep levels of contemplation. Khan et al. [24] shows that even non-expert weak LLMs can supervise expert LLMs if we allow the two LLM experts to engage in debates. Moreover, Zhao et al. [43] and Gu et al. [20] show interesting case studies where LLMs are engaged in simulated competitive environments and demonstrate human-like strategies.
Figure 2: The process of a Lincoln-Douglas-style peer-battle with the actions used. The <Think> action can be used by the candidates freely and is only visible to the candidate itself.
@@ 3 The Auto Arena Framework
The Auto-Arena framework consists of three stages: question generation, peer battles, and committee discussions. These three stages are run sequentially and fully automated with LLM-based agents. Figure 1 shows the overall design. All prompts are included in Appendix A.1.
@@ 3.1 Question Generation
For debate questions, as using a static dataset could incur data contamination concerns and result in unfair evaluations, we ask an LLM examiner agent to dynamically generate questions. The examiner agent could be any capable LLM. Similar to MT-Bench [44], the generated questions cover 8 common categories in real-life conversations: writing, roleplay, extraction, reasoning, math, coding, STEM knowledge, and humanities/social science knowledge. In the prompt, the examiner is provided with a sample question and encouraged to generate diverse and difficult questions to ensure the depth and width of the debates examined. Examples of the generated questions is shown in Appendix A.2.
Specifically, as the examiner agent will also participate in the following debates, previous methods [5] could incur self-enhancement bias as they ask the examiner agents to only devise questions that they are confident about. We try to avoid such bias with two designs: 1. We do not disclose to the examiner that it will participate in this tournament. 2. We do not ask the examiner to only generate questions that it can solve. Thus, the questions do not cause much enhancement bias for the examiner.
@@ 3.2 Peer Debate
After the questions are drafted, we conduct peer battles around these questions among the candidate LLMs. In one peer battle, two candidate LLMs (A and B) debate around the given question, point out the opponent’s weaknesses, and raise follow-up questions that the opponent may fail to respond to.
In the peer battle, each candidate LLM can use 4 types of actions: 1. Think, where the candidate thinks about the question and plans its strategy. This action is hidden to the opponent. 2. Respond, where the candidate responds to the aforementioned question. 3. Criticize, where the candidate points out the loopholes and mistakes in the opponent’s previous responses. 4. Raise, where the candidate raises follow-up questions that are specifically designed to make the opponent expose its weaknesses.
The peer battle is designed according to the Lincoln-Douglas debate format  [^3] https://en.wikipedia.org/wiki/Lincoln-Douglas_debate_format, which is a competitive one-to-one debate used for popular competitions such as National Speech and Debate Association competitions. Overall, the peer battle consists of 3 rounds, where the candidates take turns to speak. The entire dialogue history is visible to both candidates. The process is illustrated in Figure 2. In the first round, A gives an initial response to the examiner’s question; B criticizes the weaknesses in A’s response and raises a targeted follow-up question; and A responds to B’s question. The second round has the same structure, with A and B roles reversed. In the third round, A and B cross-examine each other: A starts by criticizing B’s previous loopholes and raises follow-up questions. After responding, B also criticizes A’s loopholes and raises questions. A concludes the battle by responding again. In this process, both A and B get an equal number of each action to ensure fairness. To further reduce position bias, A and B’s order is randomly determined at the beginning of each debate.
Depending on which turn it is, we provide an action guide to the candidate, specifying the objectives and corresponding actions for this turn. Similar to human debate competitions, we time the candidates by providing a maximum length, which is also specified in the prompts. Any responses beyond the required length will be cut off. This design mitigates verbosity bias in LLM-as-a-judge [44], where LLM judges prefer longer and more verbose responses.
@@ 3.3 Committee Discussions
After the peer battle takes place, a committee of LLM judge agents will collectively determine the winner. The committee is always selected as the best five LLMs (except the participants) according to the current ranking. In the first round, the committee is initialized with MMLU [22] scores to approximate LLM performances. Judges are first asked to read through the battle history, elaborate judgment reasons, and give a decision on whether A is better, or B is better, or if there is a tie.
After the initial judgments are formed, the committee engages in a discussion. Each judge reads the other judge’s verdicts, decides whether to adjust the ratings, and elaborate the reasons. The collective intelligence introduces diverse viewpoints, improves judgment quality, and mitigates single-model bias. Finally, the winning candidate is decided by the majority voting of the discussed judgments.
For logical-reasoning questions that have ground-truth answers (reasoning, code, math), LLM-as-a-judge is known to show weak performances in judging the quality of responses [44]. We adopt prior approaches to establish the reference-based judge [44]. Specifically, we ask the strongest committee member to generate a reference answer and provide it to the judge when evaluating these responses.
@@ 4.1 Experimental Setup
For the main experiment, we first select 7 best or latest models from each popular model family on the top 30 list on Chatbot Arena with more than 10k votes each. To construct a comprehensive leaderboard [^4] https://huggingface.co/spaces/Auto-Arena/Leaderboard, we further add 10 models. Appendix A.4 provides a detailed list.
@@
For the baselines, we consider widely used evaluation benchmarks, including static datasets with fixed metrics and model-based metrics. A comparison table is shown in Appendix A.5.
1. Static datasets with fixed metrics: (1) MMLU (Massive Multitask Language Understanding) [22], an extensive benchmark that covers 57 subjects and tests both world knowledge and problem-solving ability; (2) OpenLLM Leaderboard [6], a benchmark averaging performance metrics on 6 datasets.
2. Static datasets with model-based metrics: (1) MT-Bench [44], a set of 80 multi-turn questions. Model responses are graded by GPT-4; (2) MT-Bench Hard [26], a benchmark dataset with 1,000 challenging user queries collected on Chatbot Arena. Model responses are graded by GPT-4; (3) Length-Controlled AlpacaEval [19], a benchmark based on AlpacaFarm evaluation set, which tests models’ abilities to follow general user instructions. Model responses are graded by GPT-4.
@@
Among the 7 participants, we conduct a swiss-style tournament: For $n$ participants, instead of pairing each participant with $(n-1)$ others, a swiss-tournament pairs each player with $\lceil log_{2}(n)\rceil$ players of similar rankings without repeats. This design effectively reduces computational costs of ranking $n$ models from $O(n^{2})$ to $O(nlog_{2}(n))$. A cost analysis is included in Appendix A.5.
Each candidate pair engages in 40 peer battles, with 5 GPT-4-generated questions from each of the 8 categories. As each battle consists of 3 rounds (each candidate speaks for 4 times), the competition scale is approximately the same as MT-Bench (80 questions, each candidate speaks twice). In the tournament, the rating scores are calculated with the ELO rating system [4, 7], which has become the standard practice in competitive games such as chess. Similar to the Chatbot Arena score calculation procedure [11], we compute the Bradley-Terry (BT) coefficients [8] for better statistical estimation.
We initialize the Swiss tournament rankings according to MMLU scores, which is a static approximation of model performances. At the end of each pairing, we re-calculate ELO scores of current models. The committee is selected to be the remaining 5 models (besides the candidates). When the participant number increases, the committee is selected as the best 5 LLMs based on current ELO rankings. After the initial judgments, the committee members engage in one round of discussion.